# LLM-SQL-Solver: Can LLMs Determine SQL Equivalence?

### v2

@@ 4.3 Judging NL2SQL Generation
There are two notions that one can take in judging the equivalence of two SQL queries, i.e., the semantic and relaxed notions. We use the proposed Miniature & Mull and Explain & Compare prompts to evaluate the equivalence of SQL queries following semantic and relaxed notions respectively.
In Table 2, we present the breakdown of the alignment between human preference and the execution match/LLM judgement on the SQL pairs (gold SQL from Spider and generated SQL from Chat, DIN, and DAIL). We observe that the relative ranking between Chat, DIN, and DAIL are preserved for both execution matches and LLM judgements. Since the generated SQL queries in Chat are from ChatGPT with zero-shot prompting, we expect the SQL queries from DIN and DAIL to achieve better accuracy. Since DAIL applied the self-consistency techniques (resource-intensive Gao et al. (2023)), we expect DAIL to be more likely to produce the best answers. As shown in Table 2, the ranking among Chat, DIN, and DAIL based on execution accuracy, GPT 3.5 turbo judgments, and GPT 4 turbo judgements with varying prompts, all aligned with our expectation.
4.3.1 Agreement between LLMs Judgements and Execution Match
We observe that for a majority of the questions, LLMs judgements and execution matches agree with each other. In particular, as shown in, Table 3, agreements between LLMs and execution match for the semantic equivalence notion have at least 69.6%; similarly, the agreements between LLMs and execution match for the relaxed equivalence notion have at least 82.4%. We believe this is an encouraging signal that implies the good quality of execution matches and the good quality of LLMs judgements. Assuming the LLMs semantic equivalence judgement is highly accurate, a high percentage of agreements indicate the execution match correctly identified many semantically different SQL queries and the underlying database contains high quality data points (up to 1000 human hours were spent on creating the Spider Benchmark Yu et al. (2018)). If the underlying database does not contain enough high quality data points, then many of the non-semantically equivalent SQL queries will return the same results. Assuming the underlying database is well designed and the execution match is highly accurate, then a high agreement between execution match and LLMs are obviously desirable.
@@ 4.3.1 Agreement between LLMs Judgements and Execution Match
Table 3: We compare the agreement between LLMs judgements and execution matches.
@@ 4.3.2 Human Evaluation on Samples
Although, the agreements between execution accuracy and LLM’s judgments are high, there is a still lots of questions with disagreements. In this section, we provide a detailed analysis on the sampled questions from the set of questions where execution match and LLM’s judgement differs. Since there are hundreds of such questions (different results between execution match and LLM) per data source and per GPT models, we sample 10 questions randomly from the SQL query pairs where LLMs and execution match disagree.
Equivalence Evaluated Manually To evaluate the equivalence between two SQL queries, we use the human preference approach. Three of the authors manually check the equivalence of the sampled questions where LLMs and execution match disagree. In addition, we have also differentiated the equivalence between the semantic notion and relaxed notion (in essence, the former is objective and the latter is subjective). To ensure the quality of human judgement, the annotations are conducted independently, and we use the majority vote to determine the label for each pair of SQL queries (i.e., each pair of SQL queries are examined by three times independently).
@@
Table 4: We compare the human preference between LLM judge and execution match on 10 randomly sampled questions from different data sources. The questions are selected from a set where LLMs and execution match have different opinions on these questions
@@
When using the semantic equivalence notion, we find GPT 3.5 turbo may not deliver better judgment than execution match, but GPT 4 turbo showcase better alignments. When the execution match detect discrepancy in the SQL queries’ outputs, it often indicates these two SQL queries are not semantically equivalent. There are a few rare cases that we observed which the SQL queries are semantically equivalent but execution match determines not equivalent. For example, when the provided database in Spider contains column values that can not be decoded (similar issues are also raised  [^5] https://github.com/taoyds/spider/issues/46), semantically equivalent queries can be misclassified by execution match.
@@
Figure 2: The correct operation rates for LLMs. Correct operations imply the LLM correctly executed and reasoned about the SQL queries on its own provided simple database.
@@
In Table 4, GPT 3.5 turbo aligned with human experts in the range of two to five out of ten. Diving deep into the questions which GPT 3.5 turbo wrongly classified the SQL queries, we find that there are common reasons for the wrong classifications: Distinct entries vs duplicated entries. GPT 3.5 turbo often failed to consider the difference between selecting unique rows and selecting rows that may be duplicated; Subtle difference in column names. When the gold SQLs and generated SQLs share very similar SQL structure but contains subtle differences in the columns used such as group by id v.s. group by student_id, GPT 3.5 turbo may fail to notice the small difference; Hallucination. GPT 3.5 turbo may returns wrong results from executing the SQL queries on its own created simple database instance and sometimes when the outputs are indeed identical, it may still state that the outputs are different. This hallucination snowballing effect Zhang et al. (2023a) is one of the top reason why GPT 3.5 turbo perform worse compared to GPT 4 turbo.
@@
When GPT 4 turbo is used as the judge, GPT 4 turbo constantly produced more accurate or comparable results to execution match, with at least 5 out of 10 questions for DAIL and at most 10 out of 10 questions for Chat. One interesting observation is that GPT 4 turbo correctly identified the difference between order by asc/desc limit 1 v.s. where col = select min/max(col). The former will always return one row, as limit 1 is used. The latter, however, may return many rows when these rows share the same min or max value. Execution match often failed to capture this difference. Although GPT 4 turbo judgment is much better than GPT 3.5 turbo, GPT 4 turbo may still miss on the small differences such as similar but different column names (such as song_name v.s. name), overlook the fact that extra columns are selected, and also sometimes hallucinate. As shown in Figure 2, we looked into how well does LLMs executed and reasoned about the SQL queries on its own generated simple database. We find that GPT 4 turbo has much higher correct operation rates which leads much better response comapred with the weaker model.
One common phenomena that lead to false answers in execution match is due to the return of the empty tuple. When the underlying database does not contain tuples that match the predicate, gold SQLs and generated SQLs will always match, irrespective of how different they are. Another interesting observation raised from a few comparisons is about the case sensitivity. OpenAI GPT models think databases are by default case insensitive; Spider benchmark, on the other hand, uses the SQLite databases which are case sensitive by default. Many popular database are case insensitive such as SQL Server Mistry and Misner (2014) and MySQL DuBois (2013). We believe that treating case sensitivity in string comparison as an excessive step, and therefore, we operate under the assumption that the underlying database instance is case insensitive when evaluating semantic equivalence.
@@
Relaxed Equivalence When using the explain and compare prompt for determining the relaxed equivalence, both GPT 3.5 turbo and GPT 4 turbo are consistently better aligned with human judgment or at par compared to execution match, as shown in Table 4. Recall, the relaxed notion aims to illustrate the degree to which humans perceive two queries as being similar in their meaning, and mimics the opinion of the users of NL2SQL services. The positive results on the alignment between LLMs judgments with human preferences indicate optimizing for LLMs judgments accuracy is a better optimization target than execution match accuracy.
@@
The main reasons for why LLMs judgements are better aligned with human preference than execution match is that LLMs can zoom into the query logic and determine whether the two SQL queries share the same logic and contain the same truth. Consider the following example, order by col asc/desc limit 1 v.s. where col = select min/max(col). They may return different or the same outputs after execution (depends on the underlying database instance). Since execution match strictly checks the outputs, it may sometimes return equivalent and other times return not equivalent. However, LLMs can correctly identify that these two queries are selecting the rows with minimal or maximal value and these two queries are equivalent under the relaxed notion,
One interesting observation we found is about null values. GPT often consider COUNT(*), COUNT(primary key), and COUNT(DISTINCT primary key) to be different. The reason is that GPT models think that null values can be inserted into the column with primary key constraint. In fact, the SQLite databases [^6] https://www.sqlite.org/lang_createtable.html, by default, allow null values to be inserted to primarily key. However, in many database primary key constraints explicitly states that null values are not allowed [^7] https://docs.oracle.com/javadb/10.10.1.2/ref/rrefsqlj13590.html. Based on the relaxed equivalence definition, we consider records in the primary columns to be not null. Although LLMs correctly argued that when null values are presented in the primary key column, then the given SQL queries lead to different results (such as COUNT(*) considers null values as unique entries, where as COUNT(primary key) ignores rows with null values), we consider these not as a significant issue.
@@ 5 Discussion and Limitation
There are some limitations or areas of improvement to this work. First, we only evaluated the capability of Large Language Models of OpenAI. There are more other models such as Llama2 7B Touvron et al. (2023), Vicuna Chiang et al. (2023), Alpaca Taori et al. (2023a) which also showcase strong understanding and reasoning for logical tasks Chen et al. (2021); Yue et al. (2023); Yan et al. (2023b). Nevertheless, we believe that GPT 3.5 turbo and GPT 4 turbo exemplify the current status of GPT model family’s capabilities. In addition, in the manually evaluating the equivalence under two different views (semantic and relaxed), it is done over the sampled questions. Since there are more than two thousand of SQL queries per data source and equivalence view, manually checking all of them are arguably infeasible. In fact, even checking a pair of SQL queries for semantic view takes around 10 minutes for experts in SQL. Although the questions are sampled, we observe a consistent trends between LLMs’ judgements and execution matches (such as GPT 4 turbo’s responses are consistently better aligned with human preferences).
@@ 6 Conclusion
In this paper, we studies and answer two main questions: i) Can current LLMs be used to help data engineers in checking semantic equivalences between SQL queries and ii) should we use LLMs to judge the quality of generated SQL queries for text-to-SQL benchmarks. Through extensive evaluations and manually examining the quality of the answers based on the alignment with human preferences, we find current LLMs (especially GPT 4 turbo), using our proposed prompt techniques, can assists engineers in determining semantic equivalent queries, but may yield incorrect responses. On the other hand, when comparing the generated SQL query with the gold SQL query with relaxed notion of equivalence, LLMs’ responses constantly achieve high alignments with human preferences. Moreover, another advantage of using LLMs to determine SQL equivalence is that it can avoid the construction of sophisticated databases, and no data stored need to be accessed. To the best of our knowledge, this is the first work studying the ability of LLMs in determining SQL equivalence. As the learning ability in LLMs continuously to advance, we believe that using these models to check SQL equivalence is a promising direction.

### v3

@@ 4.3 Judging NL2SQL Generation
There are two notions that one can use in judging the equivalence of two SQL queries, i.e., the semantic and relaxed notions. We use the proposed Miniature & Mull and Explain & Compare prompts to evaluate the equivalence of SQL queries following semantic and relaxed notions respectively.
In Table 2, we present the breakdown of the alignment between human preference and the execution match/LLM judgement on the SQL pairs (gold SQL from Spider and generated SQL from Chat, DIN, and DAIL). We observe that the relative ranking among Chat, DIN, and DAIL are preserved for both EX and LLM judgements. Since the generated SQL queries in Chat are from ChatGPT with zero-shot prompting, we expect the SQL queries from DIN and DAIL to achieve better accuracy. Since DAIL applied the self-consistency techniques (resource-intensive Gao et al. (2023)), we expect DAIL to be more likely to produce the best answers. As shown in Table 2, the ranking among Chat, DIN, and DAIL based on execution accuracy, GPT 3.5 turbo judgments, and GPT 4 turbo judgements with varying prompts, all aligned with our expectation.
4.3.1 Agreement between LLMs’ Judgements and Execution Accuracy
For a majority of the questions, LLMs’ judgements and EX agree with each other. In particular, as shown in, Table 3, agreements between LLMs and EX for the semantic equivalence notion have at least 69.6%; similarly, the agreements between LLMs and EX for the relaxed equivalence notion have at least 82.4%. We believe this is an encouraging signal that implies the good quality of the Spider benchmark and LLMs’ judgements. The Spider benchmark contains high quality data points (1000 human hours spent Yu et al. (2018)). Creating these high quality databases, which is necessary for good EX, are expensive and does not scale. Assuming the underlying database is well designed and the EX is highly accurate, then a high agreement between EX and LLMs are obviously desirable.
@@ 4.3.1 Agreement between LLMs’ Judgements and Execution Accuracy
Table 3: We compare the agreement between LLMs’ judgements and execution matches.
@@ 4.3.2 Human Evaluation on Samples
Although, the agreements between execution accuracy and LLM’s judgments are high, there are still lots of questions with disagreements. In this section, we provide a detailed analysis on the sampled questions from the set of questions where EX and LLM’s judgement differs. Since there are hundreds of such questions (different results between EX and LLM) per data source and per GPT models, we sample 10 questions randomly from the SQL query pairs where LLMs and EX disagree.
Equivalence Evaluated Manually To evaluate the equivalence between two SQL queries, we use the human preference approach. Three of the authors manually checked the equivalence of the sampled questions where LLMs and EX disagree. In addition, we have also differentiated the equivalence between the semantic notion and relaxed notion (in essence, the former is objective and the latter is subjective). To ensure the quality of human judgement, the annotations are conducted independently, and we use the majority vote to determine the label for each pair of SQL queries.
@@
Table 4: We compare the human preference between LLM judge and EX on 10 randomly sampled questions from different data sources.
@@
When using the semantic equivalence notion, we find GPT 3.5 turbo may not deliver better judgment than EX, but GPT 4 turbo results in better alignments. When EX detects discrepancy in the SQL queries’ outputs, it often indicates these two SQL queries are not semantically equivalent. There are a few rare cases that we observed in which the SQL queries are semantically equivalent but EX determines not equivalent. For example, when the provided database in Spider contains column values that can not be decoded (similar issues are also raised sythello , semantically equivalent queries can be misclassified by EX.
@@
Figure 2: The rates of LLM correctly executed the SQL queries on its own provided simple database.
@@
In Table 4, GPT 3.5 turbo aligned with human experts in the range of two to five out of ten. Diving deeper into the questions which GPT 3.5 turbo wrongly classified the SQL queries, we find that there are common reasons for the wrong classifications: Distinct entries vs duplicated entries. GPT 3.5 turbo often failed to consider the difference between selecting unique rows and selecting rows that may be duplicated; Subtle difference in column names. When the gold SQLs and generated SQLs contain only subtle differences in the columns used such as group by id v.s. group by student_id, GPT 3.5 turbo may fail to notice these differences; Hallucination. GPT 3.5 turbo may return wrong results when executing the SQL queries on its own created simple database instance and sometimes when the outputs are indeed identical, it may still state that the outputs are different. This hallucination snowballing effect Zhang et al. (2023a) is one of the top reason why GPT 3.5 turbo performs worse compared to GPT 4 turbo.
@@
When GPT 4 turbo is used as the judge, it constantly produced more accurate or comparable results to EX, with at least 5 out of 10 questions for DAIL and at most 10 out of 10 questions for Chat. One interesting observation is that GPT 4 turbo correctly identified the difference between order by asc/desc limit 1 v.s. where col = select min/max(col). The former will always return one row, as limit 1 is used. The latter, however, may return many rows when these rows share the same min or max value. EX often failed to capture this difference. Although GPT 4 turbo’s judgment is much better, GPT 4 turbo may still miss some minor differences such as similar but different column names (such as song_name v.s. name), overlook the fact that extra columns are selected, and also sometimes hallucinates. As shown in Figure 2, we explored how well do LLMs execute and reason about the SQL queries on their own generated simple database. GPT 4 turbo has much higher correct operation rates which leads to much better responses compared with the weaker model. One common phenomena that lead to false answers in EX is due to the return of the empty tuple. When the underlying database does not contain tuples that match the predicate, gold SQLs and generated SQLs will always match, irrespective of how different they are. Another interesting observation raised from a few comparisons is about the case sensitivity. OpenAI GPT models consider databases are by default case insensitive; Spider benchmark, on the other hand, uses the SQLite databases which are case sensitive by default. However, many popular database are by default case insensitive such as SQL Server Mistry and Misner (2014) and MySQL DuBois (2013). Therefore, we operate under the assumption that the underlying database instance is case-insensitive when evaluating semantic equivalence.
@@
Relaxed Equivalence. When using the Explain & Compare prompt for determining the relaxed equivalence, LLMs are consistently better aligned with human judgments than EX, as shown in Table 4. The relaxed notion aims to illustrate the degree to which humans perceive two queries as being similar in their meaning, and mimics the opinions of users. The positive results on the alignment between LLMs’ judgments with human preferences indicate optimizing accuracy for LLMs’ judgments is a better optimization target than EX.
@@
The main reasons for why LLMs’ judgements are better aligned with human preference than EX is that LLMs can zoom into the query logic and determine whether the two SQL queries share the same logic and contain the same truth. Consider the following example, order by col asc/desc limit 1 v.s. where col = select min/max(col). LLMs can correctly identify that these two queries are selecting the rows with minimal or maximal value. Since EX strictly checks the outputs, it may sometimes return equivalent and other times return not equivalent. One interesting observation we found is about null values. GPT often consider COUNT(*), COUNT(primary key), and COUNT(DISTINCT primary key) to be different. The reason is that GPT models consider that null values can be inserted into a column with primary key constraint. In fact, the SQLite databases, by default, allow null values to be inserted to primary key column. However, in many database primary key constraints explicitly states that null values are not allowed Oracle . Based on the relaxed equivalence definition, we consider records in the primary columns to be not null. Although LLMs correctly consider that having null values in the primary key column may lead to different results (note that COUNT(*) considers null values as unique entries, where as COUNT(primary key) ignores rows with null values), we consider these not as a significant issue.
@@ 5 Discussion and Limitation
In this discussion section, we would like to highlight the broader implications and potential integration of our proposed LLM-SQL-Solver. During the SQL rewrite phase of the query optimization engine, the input SQL query is transformed into a new SQL query that can be optimized more easily. The key challenge in query rewriting is to determine if the input query and the rewritten query are indeed semantic equivalent. The LLM-SQL-Solver can be directly used as a tool, during the rewriting phase, to check semantic equivalence and ensure bugs (e.g., COUNT bug Ganski and Wong (1987)) do not occur in the data system. Furthermore, the integration of the LLM-SQL-Solver as a text-to-SQL evaluation metrics using the relaxed equivalence notion could improve the overall quality of user experience and enable more human-like capabilities.
Concurrent with our research, the chain-of-table approach Wang et al. (2024) independently explores a similar concept, allowing LLMs to form a reasoning path through table modifications to improve question answering from tabular data, akin to our Miniature & Mull prompt. However, while our technique focuses on determining query semantic equivalence through table creation and modification, their goal is to deepen the understanding of the tabular data itself.
There are some limitations or areas of improvement to this work. First, we only evaluated the capabilities of the Large Language Models of OpenAI. There are other alternative models such as Llama2 7B Touvron et al. (2023), Vicuna Chiang et al. (2023), Alpaca Taori et al. (2023a) which also showcase strong understanding and reasoning for logical tasks Chen et al. (2021); Yue et al. (2023); Yan et al. (2023b). Nevertheless, we believe that GPT 3.5 turbo and GPT 4 turbo exemplify the current status of the GPT model family’s capabilities. In addition, our evaluations on Spider Benchmark were conducted over randomly sampled questions, in which we have to manually label the equivalence of the two different notions (semantic and relaxed). Since there are nearly ten thousand pairs of SQL queries, manually checking all of them are arguably infeasible. In fact, even checking a pair of SQL queries for semantic view takes around 10 minutes for experts in SQL. Although the questions are sampled, we observe a consistent trends between LLMs’ judgements and execution matches (such as GPT 4 turbo’s responses are consistently better aligned with human preferences).
@@ 6 Conclusion
In this paper, we studied and answered two main questions: i) Can current LLMs be used to help engineers in checking semantic equivalences between SQL queries and ii) should we use LLMs to judge the quality of generated SQL queries for text-to-SQL benchmarks. To the best of our knowledge, this is the first work studying the ability of LLMs in understanding SQL queries and determining SQL equivalences. Through extensive evaluations and manually examining the quality of the answers based on the alignment with human preferences, we find current LLMs (especially GPT 4 turbo), using our proposed prompt techniques, can assists engineers in determining semantic equivalent queries, but may yield incorrect responses. On the other hand, when comparing the generated SQL query with the gold SQL query with our relaxed notion of equivalence, LLMs’ responses constantly achieve high alignments with human preferences. Moreover, our proposed LLM-SQL-Solver avoids accessing stored user data, and hence no sophisticated databases need to be constructed.